# Is it Possible to Edit Large Language Models Robustly?

### v1

@@ 3.2 Edited LLM
53.9
26.2
55.9
20.8
20.9
19.7
40.3
12.3
78.0
6.3
93.9
4.9
 + Fill-in-the-blank
26.5
30.7
40.3
23.0
12.5
16.8
22.9
14.1
2.9
18.6
58.7
13.4
 + Reference resolution
19.5
35.6
26.1
29.5
8.7
15.1
15.1
12.5
18.9
6.2
72.3
5.5
@@
58.7
21.8
55.4
19.0
20.1
18.0
33.5
13.0
20.5
2.5
73.5
10.3
 + Fill-in-the-blank
26.7
30.8
39.1
22.7
12.5
16.4
20.3
13.8
2.5
17.8
33.0
18.2
 + Reference resolution
20.7
30.7
25.7
26.0
6.6
13.5
11.9
11.7
9.5
2.0
50.6
9.2
[…]
54.2
26.0
51.8
17.2
15.1
0.8
31.0
1.6
70.5
4.7
92.0
4.2
 + Fill-in-the-blank
31.4
30.0
44.0
22.1
13.1
14.5
22.2
11.3
2.3
17.2
61.4
13.1
 + Reference resolution
23.4
28.1
29.0
20.7
9.5
0.9
16.0
1.2
24.5
5.7
58.1
4.3
[…]
55.8
21.0
51.8
16.1
16.0
0.8
30.6
1.6
29.3
3.6
78.4
5.5
 + Fill-in-the-blank
31.3
28.8
43.0
20.8
13.0
13.2
21.7
10.7
2.1
15.6
46.7
13.9
 + Reference resolution
23.0
24.6
27.0
18.8
10.1
0.7
17.0
0.8
15.5
5.3
45.3
3.6
[…]
44.8
42.9
58.7
39.1
40.1
37.8
47.3
35.2
20.0
46.3
7.4
47.4
Table 1: Results on CounterFact and zsRE with Llama-7b and 13B foundation models. acc: accuracy, rev: reversion.
(d) Raising doubts. In realistic deployment, a special case is when the user questions the statement about the target knowledge or even doubts the factuality. This situation cannot be ignored for practical use, i.e., the successfully edited knowledge memory needs to be robust when questioned. Two prompts of challenging are used in our experiments. One is only to doubt the target knowledge, and another expresses an explicit negative objection to the output and asks about the original answer $o$ (Appendix B.3).
To sum up, six groups of attacking prompts are prepared. First is (i) fill-in-the-blank clozes. Then we have (ii) long context with related knowledge, (iii) noisy long context, (iv) simulated dialogue, and (v) noisy dialogue, which are all followed by three kinds of prompts, i.e., plain prompts, plain prompts with reference, and fill-in-the-blank questions. The last is the questions that (vi) raise doubts. Section 4 will present the empirical studies on the proposed attacking prompts.
4 Experiments
This section describes our empirical studies on the proposed attacking strategies. The setups are presented followed by results and analysis.
4.1 Dataset
Two mainstream datasets are used for experiments.
CounterFact Meng et al. (2022) is proposed for evaluating significant counterfactual edits. Each sample is annotated with the explicit $(s,r,o)$ triplet, the new object $o^{\prime}$, and rephrased prompts. The plain prompt is fixed templates according to $r$, filled with $s$. The rephrased prompts are semantically equivalent expressions.
zsRE De Cao et al. (2021); Levy et al. (2017), zero-shot relation extraction, derives from a factual question-answering task. Following existing work Yao et al. (2023), the alternative answer is used as the new answer $o^{\prime}$. Each sample includes the subject $s$, the true answer $o$, the new answer $o^{\prime}$, the question, and its rephrased version.
4.2 Baselines and Implementation
The experiments focus on popular editing methods of different types, including (i) locate-then-edit methods: KN Dai et al. (2022b), ROME Meng et al. (2022), MEMIT Meng et al. (2023); (ii) external module-based methods: SERAC Mitchell et al. (2022b) relies on an external memory, while MEND Mitchell et al. (2022a) works with a hypernetwork. (iii) prompt-based method: IKE Zheng et al. (2023). Llama-2-7B and 13B-chat Touvron et al. (2023) are adopted as the foundation models. Details setups are presented in Appendix B.4.
(a) Frequency
(b) Connection
(c) Co-occurence
@@ 4.2 Baselines and Implementation
Metrics. All metrics are computed based on generated text from the edited model. After editing, the prompts are inputted and the model outputs are collected. Then the output is normalized. The edit is counted as a success if the new answer $o^{\prime}$ appears in the output, denoted as accuracy. We additionally compute the appearance of the original answer $o$, denoted as reversion.
4.3 Results and Analysis
Table 1 shows the main results, indicating that all of the editing methods can be vulnerable and not ready for practical use. Following are our key findings.
(i) Locate-then-edit methods and external module-based methods show differential performance, while the prompt-based method is better suited for the LLM. Concretely, ROME, MEMIT, SERAC, and IKE achieve a nearly perfect score on the plain prompt. KN loses the effectiveness and MEND shows a significant preference for original answers $o$ on a comparable level of the model before the edit. However, the methods with promising scores can fail to face our attacks.
(ii) ROME and MEMIT show relatively subside decreases in attacks including format changes, related or noisy knowledge probing, and dialogue history. But they suffer from the reference of subjects and compromise to doubtful questions.
(iii) The performance of SERAC almost totally relies on the scope classifier. Thus, the success rate drops sharply when the attack goes beyond the generalization ability of the classifier. Although the long inputs are truncated from the left side, the change of question form can still easily bypass the classification.
(iv) The prompt-based approach, IKE, generally achieves the best robustness. This indicates that proper prompts leverage the instruct-following potential of LLMs to control the output. However, this can be easily attacked by input, as the user can inject any knowledge into the prompt.
(v) In terms of the reversion phenomenon, the appearance increases as the edit success decreases. Related contexts with neighbor knowledge facilitate the reversion. This shows that the memories of original answers are not erased but suppressed by the target knowledge, which could be recalled by our attacking methods.
(a) Perplexity distributions by Llama  -2-7B-chat.
(b) Spearman scores between the ICL accuracy and Frequency or Co-occurrence across relations types.
@@ 4.3 Results and Analysis
5 Knowledge Popularity Affecting Editing Robustness
Besides the extrinsic effects like various inputs, intrinsic features of involved knowledge can influence editing. This section studies Q3: Which knowledge features are correlated with the performance and robustness of editing? Based on previous results, we summarized the possible features as popularity. We first define three measurements, then proceed to analysis and findings.
5.1 Popularity Measurements
We measure the knowledge features of realistic popularity from three aspects below (Appendix C).
(i) Frequency. The frequency of an entity can be measured by how often its Wikipedia entry is visited Mallen et al. (2023). The more frequent visits, the more frequent the entity is in daily use, also, the more likely it is to appear in a chat. We use the monthly view numbers of the subject.
@@ 5.2 Analysis
(i) Existing benchmarks edit less popular entities on the aspects of Frequency, Connection, and Co-occurrence. Figure 4 shows frequencies of the entities in four datasets, including two editing benchmarks, CounterFact and zsRE, and widely accepted knowledge-intensive question-answering datasets, TriviaQA Joshi et al. (2017) and Natural Question Kwiatkowski et al. (2019). It can be observed that CounterFact and zsRE contain more entities whose monthly visits are around $10^{2}$ and $10^{3}$, while TriviaQA contains more entities that are viewed around $10^{4}$-$10^{5}$ times. Both the connection and co-occurrence show long-tail shapes. But TriviaQA decreases in a relatively gentle trend, containing more samples with higher Connection and Co-occurrence. This indicates that entities and knowledge in editing benchmarks are much less likely to show up in a realistic conversation.
@@
(ii) Language models have weaker memory for less popular knowledge, thus simplifying the editing and resulting in biased findings. We try to probe knowledge memorization by comparing the perplexities of the answers. The perplexities are computed of $o$ and $o^{\prime}$ as completions of the plain prompt on Llama. Figure 4 presents the distribution of the logarithmic perplexities difference of $o$ and $o^{\prime}$. There are 16.22% samples in CounterFact and 43.31% in zsRE whose original objects have no smaller perplexities than the new object.
@@
We also directly prompt LLMs without editing to see whether the model can recall the knowledge. Two settings are considered. (a) The plain prompt is input and the original object $o$ is expected as the completion. (b) The input follows the format of in-context learning (ICL) Brown et al. (2020), i.e., a concatenation of “Instruction, Demonstrations, Question.” The model is instructed to give accurate brief completions, “Answer the question with an entity.” This stimulates the potential of the parametric memories to the maximum extent.
Model
Llama-2-7B-chat
GPT-j
GPT-2XL
CounterFact
31.8/1.1
29.5/1.2
18.2/0.6
 w/ ICL
57.0/2.4
47.9/2.8
34.5/4.2
zsRE
20.9/4.3
7.1/3.3
@@
Table 2: Accuracy of recalling the knowledge, $o$ or $o\prime$, by the models without editing.
@@
Table 2 shows the scores on our base model, Llama-7B, and common baselines Meng et al. (2023); Yao et al. (2023), GPT-J Wang (2021) and GPT-2XL Radford et al. (2019). The plain prompt leads to diverse completions without constraints. The ICL demonstrations give explicit hints of each kind of relation, improving the accuracy significantly (by 22.7% on Llama, 18.4% on GPT-j, and 15.3% on GPT-2XL). However, there is still around half of the knowledge that can not be recalled. This indicates that in the first place, a considerable part of the knowledge to edit is not memorized with high confidence or can not be used effectively. Thus, the knowledge in benchmarks only needs injection with weak prior knowledge, possibly facing less resistance and risk of side effects. The difficulty of model editing can still be underestimated.
The correlation between knowledge popularity and parametric memory can be verified by the Spearman scores shown in Figure 4. The scores are computed between ICL accuracy and Frequency or Connection on CounterFact. Most relation types have scores around 0.1--0.3. A few relation types are negative outliers. For example, the relation [X] and [Y] are twin cities. rarely exist in memories and get various outputs. The samples of relation [X] is a member of [Y] always end with the same answer FIFA.
@@
(iii) Editing more popular knowledge is more difficult and more vulnerable to rephrasing. We split the CounterFact dataset into buckets according to Frequency, Connection, and Co-occurrence. ROME and MEMIT are applied to edit the knowledge and evaluated on the plain prompts and rephrased prompts from the original benchmark. The results are shown in Figure 5. The success on plain prompts keeps high scores and gentle decreases on the three measurements. Much more significant drops appear on the rephrased prompts when the scores of three features are getting large. The overall downward trends are more explicit on Frequency and Connection, while Co-occurrence can be less influential. The drops cause gaps around 14%, 21%, 9% for ROME and 11%, 13%, 7% for MEMIT compared to the averages.
@@
To sum up, knowledge with higher popularity has more valid parametric memory and higher portions in practical use. For LLMs, those pieces of knowledge are easier to recall and harder to effectively change by existing editing methods.
@@ 6 Conclusion
This paper systematically studies recent model editing methods under the situation of practical use and raises concerns about their robustness. We first show that confusion and hallucination occur in realistic dialogues with edited LLMs. Besides the user-AI interaction, we also rephrase the prompts by changing the format or adding context to attack editing. The vulnerability of target knowledge is shown in our experiments. For more analysis, three knowledge popularity measurements are proposed. We show that more popular knowledge is memorized better, easier to recall, and harder to robustly edit for LLMs. Although editing methods show inspiring success in manipulating the memory and behaviors of LLMs, they can be problematic in practical situations.
@@
We acknowledge the limitations of this work. (i) Baseline coverage. Although this paper has considered a wide range of popular baselines, it is hard to cover all existing work. This paper involves mainstream LLM editing of different types following recent work Yao et al. (2023); Zhong et al. (2023); Zheng et al. (2023). Due to the resource limitations, this paper selects Llama-2-chat in the size of 7B, 13B as foundation models and leaves the larger size as future work. To make the attacking methods easy to automate and illustrate, we use counterfactual datasets as the benchmarks. We also consider the situation of changing a wrong memory to the correct knowledge by editing, where confusion and hallucination still exist as shown in Figure 7. (ii) Human evaluation. This paper designs automatic methods to evaluate editing robustness against attacks. However, humans can give more sophisticated attacking prompts, e.g., by asking humans to have a chat with edited models instead of GPT-4.
Appendix A Details of Edited Communicative AI
A.1 Automatic Evaluation
Figure 6 shows a human evaluation case where the edited model gets confused. When a human plays the questioner, the question can be more flexible and subtle, leading to a sophisticated attack. However, we turn to GPT-4 to design a human-free process for the convenience of experiments. The prompts input to GPT-4 is as follows.

### v2

@@ 5.1 Method
Cloze 31.3 28.8 43.0 20.8 13.0 13.2 21.7 10.7 2.1 15.6 46.7 13.9
w/ Reference 23.0 24.6 27.0 18.8 10.1 0.7 17.0 0.8 15.5 5.3 45.3 3.6
@@
Table 1: Results on CounterFact and zsRE with Llama-7b and 13B models (acc: accuracy, rev: reversion). The Direct prompt and Equivalent prompt are from benchmarks. N/A means we add no context in front of the query.
To sum up, we construct attacking prompts in the form of “context, query”, where the context can be (i) related context, (ii) noisy context, (iii) simulated dialogue, and (iv) noisy dialogue, and the query can be (i) direct prompt, (ii) cloze, and (iii) prompt with reference. We also prepare prompts that raise doubt. Section 5.2 will present results on these attacking prompts.
5.2 Experiments for RQ2
5.2.1 Datasets
Our evaluation adopts three mainstream datasets: (i) CounterFact (Meng et al., 2022) includes significant counterfactual edits. Each sample is annotated as $(s,r,o)$ triplet with a target object $o^{\prime}$. The direct prompts $x$ are fixed templates based on $r$, with their equivalent expressions $x_{e}$ also provided. (ii) zsRE (De Cao et al., 2021; Levy et al., 2017), zero-shot relation extraction, derives from a factual question-answering task. Following Yao et al. (2023), the alternative answer is used as $o^{\prime}$. Each sample is annotated as $(s,o,o^{\prime},x,x_{e})$, where $x$ and $x_{e}$ are questions. (iii) A time-changing dataset, MQUAKE-T (Zhong et al., 2023), is also incorporated to validate of our findings (Appendix C).
5.2.2 Baselines and Implementation
The experiments cover popular editing methods of different types, including (i) locate-then-edit methods: KN (Dai et al., 2022b), ROME (Meng et al., 2022), MEMIT (Meng et al., 2023); (ii) external module-based methods: SERAC (Mitchell et al., 2022b) relies on an external memory, while MEND (Mitchell et al., 2022a) works with a hypernetwork. (iii) prompt-based method: IKE (Zheng et al., 2023). Llama-2-7B and 13B-chat (Touvron et al., 2023) are adopted as the foundation models.
Metrics. All metrics are computed based on auto-regressively generated texts from the edited models. The test is considered successful if the new answer $o^{\prime}$ appears in the normalized output, with the proportion referred to as accuracy, dubbed as acc. We also compute the appearance of the original answer $o$, reversion, dubbed as rev. Detailed settings are presented in Appendix B.4.
5.3 Analysis for RQ2
Table 1 indicates that popular editing methods exhibit vulnerabilities and are not yet ready for practical use. Key findings are presented as follows.
(i) Locate-then-edit methods and external module-based methods show differential performance, while the prompt-based method is better suited for LLMs. Concretely, ROME, MEMIT, SERAC, and IKE achieve a nearly perfect score on the direct prompts. KN almost loses its effectiveness. MEND achieves a success rate of around half. However, the methods with promising scores can fail to face our attacks.
(ii) ROME and MEMIT show relatively moderate decreases in attacks of lengthy contexts but suffer from query changes (cloze form and reference resolution) and doubting questions. Their performance also decreases on the larger-size model.
(iii) The performance of SERAC mostly relies on the scope classifier. Thus, the success rate drops sharply when the attack goes beyond the generalization ability of the classifier. Although the long inputs are truncated from the left side, the cloze format can still bypass the classification. This indicates enormous potential for SERAC by classifier improvement, as the performance could match IKE if we assume a perfect classifier.
(iv) The prompt-based approach, IKE, generally achieves better robustness, showing that in-context learning (Brown et al., 2020) stimulates the generalization and instruct-following of LLMs to control the output. However, the performance depends on demonstrations, which can be compromised in practical interactions, as the user can inject knowledge into the input. When the edit is unknown, the retrieved demonstrations can be a sub-optimal set.
(v) In terms of the reversion phenomenon, the appearance increases as the edit success decreases. Long contexts with neighbor knowledge largely increase the reversion. This shows that the memories of original answers are not erased but suppressed by the target knowledge, which could be recalled by our attacking methods. [^1] Appendix B.5 provides a fine-tuning baseline.
@@ 5.3 Analysis for RQ2
6 RQ3: Knowledge Popularity Affecting Editing Robustness
Besides the extrinsic effects like various inputs, this section studies RQ3, the influence of intrinsic knowledge features on editing, especially the popularity.
6.1 Method
We define the knowledge popularity and its measurements from three aspects (Appendix D).
(i) Frequency. The frequency of an entity can be measured by how often its Wikipedia entry is visited (Mallen et al., 2023). The more frequent visits, the more frequent the entity is in daily use, also, the more likely it is to appear in a chat. We use the monthly view number of the subject.
@@ 6.2 Analysis for RQ3
(i) Existing benchmarks edit less popular knowledge on the aspects of Frequency, Connection, and Co-occurrence. Figure 3 shows frequencies of the entities in four datasets, including two editing benchmarks, CounterFact and zsRE, and three widely accepted knowledge-intensive question-answering datasets, TriviaQA (Joshi et al., 2017) and Natural Question (Kwiatkowski et al., 2019). It can be observed that editing benchmarks contain more entities with Frequencies around $10^{2}$-$10^{3}$, while QA datasets contain more entities viewed around $10^{4}$-$10^{5}$ times. Both the Connection and Co-occurrence also decrease in slower trends in QA datasets. This indicates that entities and knowledge in editing benchmarks are much less likely to appear in a realistic conversation.
@@
(ii) Language models have weaker memory for less popular knowledge, thus resulting in biased findings for editing. We probe knowledge memorization by comparing the perplexities of the answers. The perplexities are computed of $o$ and $o^{\prime}$ as completions of the direct prompt on Llama. Figure 9 presents the distribution of the logarithmic perplexities difference of $o$ and $o^{\prime}$. There are 16.22% samples in CounterFact and 43.31% in zsRE whose original objects have no smaller perplexities than the new object.
@@
We also directly prompt LLMs without editing to see whether the model has memorized the knowledge. Two settings are considered: (a) The direct prompt is input and the original answer $o$ is expected as the completion. (b) The input follows the format of in-context learning (ICL), i.e., a concatenation of “instruction, demonstrations, direct prompt.” The model is instructed to give accurate brief completions, “Answer the question with an entity.” ICL stimulates the potential of the parametric memories to the maximum extent.
Model Llama-2-7B-chat GPT-j GPT-2XL
CounterFact 31.8/1.1 29.5/1.2 18.2/0.6
 w/ ICL 57.0/2.4 47.9/2.8 34.5/4.2
zsRE 20.9/4.3 7.1/3.3
@@
Table 2: Accuracy of probing parametric knowledge, $o$ or $o\prime$, by the models without editing.
@@
Table 2 shows the scores on our base model, Llama-2-7B-chat, and common baselines (Meng et al., 2023; Yao et al., 2023), GPT-J (Wang, 2021) and GPT-2XL (Radford et al., 2019). The direct prompt leads to diverse completions without constraints. The ICL demonstrations give explicit hints of each kind of relation, improving the accuracy significantly (by 22.7% on Llama, 18.4% on GPT-j, and 15.3% on GPT-2XL). However, about half of the knowledge still cannot be recalled. This suggests that, in the first place, a considerable portion of the knowledge to be edited is either not memorized with high confidence or cannot be used effectively." Knowledge with weak prior memory possibly has less resistance and risk of side effects. Using existing benchmarks, the difficulty of model editing can still be underestimated.
Figure 9 shows the Spearman score to verify the correlation between knowledge popularity and parametric memory (ICL accuracy). Most relation types have scores around 0.1--0.3.
@@
(iii) Editing more popular knowledge is more vulnerable to rephrasing. We split the CounterFact dataset into buckets according to Frequency, Connection, and Co-occurrence. ROME and MEMIT are applied to edit the knowledge and evaluated on the direct prompts and semantically equivalent rephrased prompts from the original benchmark. The results are shown in Figure 4. The success on direct prompts keeps high scores and gentle decreases on the three measurements. Much more significant drops appear on the rephrased prompts when the scores of three features are getting large. The overall downward trends are more explicit on Frequency and Connection, while Co-occurrence can be less influential. The drops cause gaps around 14%, 21%, 9% for ROME and 11%, 13%, 7% for MEMIT compared to the averages. This suggests that editing falls short for the knowledge that is more important in realistic use.
@@
In summary, knowledge with higher popularity tends to have more reliable parametric memory for practical use based on Frequency, Connection, and Co-occurrence. For LLMs, those pieces of knowledge are easier to recall and harder to modify by existing editing methods robustly.
7 Potential Mitigation
Our work could suggest promising directions for improving the editing robustness as follows.
(i) From the data perspective, one solution is to consider more complex inputs in the editing phase. Existing methods incorporate mechanisms for generalization to some extent (e.g., prefix sampling in ROME). We can further enhance the diversity and complexity. (ii) From the LLM-ability perspective, another solution is to develop effective pipelines integrating disentangling and reasoning workflow Khattab et al. (2022); Chern et al. (2023), e.g., to disentangle required knowledge from lengthy inputs by claim extraction or query rewriting, and then bootstrap the required (edited or original) knowledge. (iii) From the method-specific perspective, it is feasible to design targeted and lightweight approaches tailored to a certain editing method, given that the vulnerabilities of different algorithms vary based on their intrinsic problems. For instance, we can resolve references to subjects in ROME and MEMIT or detect doubtful questions in IKE. We conduct experimental validation for those mitigation strategies, each of which leads to average improvements. Please see Appendix E.
@@ 8 Conclusion
This paper systematically studies recent model editing methods under the situation of practical use and raises concerns about their robustness. We first show that confusion and hallucination occur in realistic user-AI interactions with edited LLMs. Besides, we rephrase the prompts by adding context and changing the question format to attack editing, demonstrating the vulnerability of target knowledge. For more analysis, we propose three knowledge popularity measurements and show that popular knowledge is memorized better, easier to recall, and harder to robustly edit for LLMs.
Editing methods have shown impressive success, while they can be problematic in practical situations because of existing robustness deficiencies. More importantly, this paper calls for effort on this inspiring research topic and underscores the collective focus on improving editing robustness for further application.
@@
We acknowledge the limitations of this work. (i) Coverage. Although it is hard to cover all application settings due to the resource limitations, this paper considers setups for baselines as much as possible, compared to recent work (Yao et al., 2023; Zhong et al., 2023; Zheng et al., 2023). This paper covers a wide range of mainstream LLM editing methods of different types. Llama-2 in 7B and 13B are adopted to represent the mainstream decoder-only LLM architecture. They show remarkable emergent abilities and have significant impacts as communitive AI in the open-source LLM community. We mainly consider two mainstream benchmarks for easier automation and comparison with previous works. (ii) Human evaluation. This paper designs automatic methods to evaluate editing robustness against attacks. However, humans can give more sophisticated attacking prompts and aggravate the confusion and hallucinations, e.g., by asking humans to have a chat with edited models instead of GPT-4.
Future work.
While our paper highlights concerns regarding the robustness of model editing, we also view model editing as an exciting field rich with research opportunities. We propose several directions for future work.
(i) Robustness augmentation and side effect elimination. We propose suggestions for potential mitigation methods and experiments in Section 7. The fundamental solutions for robust editing remain to be explored, including more complex scenarios like multiple edits and general capability retention Wang et al. (2024b); Ma et al. (2024).
(ii) LLM safety. In the context of information security, our findings strongly relate to the CIA triad (Confidentiality, Integrity, and Availability). Reversion poses a risk to confidentiality, as it may expose knowledge that should remain undisclosed. Hallucination threatens integrity and availability by potentially leading to the dissemination and exploitation of inaccurate content. Developing effective alignment approaches is crucial for implementing robust defense methods Patil et al. ; Suau et al. (2024).
(iii) Other applications. While we demonstrate temporary deficiencies in editing robustness for generative AI, existing editing technologies have demonstrated reasonable performance in terms of success rates and locality. These technologies have potential applications in scenarios requiring strict limitations or precise triggering conditions, such as backdoor or authorization systems Li et al. ; Qiu et al. (2024). The application scope can extend far beyond knowledge-related cases, encompassing a wide range of uses in AI systems.
In essence, this paper calls for effort on this inspiring research topic and underscores the collective focus on enhancing editing robustness for reliable practical application.
Appendix A Details of Experiments for RQ1
A.1 Evaluation for RQ1
A.1.1 Automatic Dialogue Generation
Chatting to the edited communitive AI is an online task, which means the “user” must receive the agents’ output and give real-time feedback in the next utterance. For the convenience of experiments, we turn to GPT-4 to design a human-free process. The prompts input to GPT-4 is as follows. With an example and those instructions, we find that GPT-4 can generate questions highly related to the target knowledge in various expressions (mostly within a 2-hop neighbor knowledge). The dialogues are limited to 5 turns.